On 1 June 2023 Safe Boards Ltd invested in five hundred 7 per cent, ten-year Teleco bonds with a face value of $100 each. The bonds were issued at face value. On 30 June 2023, the Teleco bonds, which are traded in an active market, had a market value of $105. Answer each part independently.

a) State whether Safe Boards Ltd can classify the Teleco bonds as being measured at amortised cost. If measured at amortised cost, give the amount at which the bonds should be reported in the statement of financial position at 30 June 2023.

b) If the bonds were acquired for speculative purposes, give the amount at which the bonds should be reported in the statement of financial position at 30 June 2023. If a change in fair value is recognised, where should it be recognised?
Answer #1

Answer to Part a :

Yes, Safe Boards Ltd can classify the Teleco bonds as being measured at amortised cost. As per IFRS 9 : Financial Instruments, a financial asset is measured at amortised cost if both of the following conditions are met :

1. The asset is held within a business model whose objective is to hold assets in order to collect contractual cash flows; and

2. the contractual terms of the financial asset give rise on specified dates to cash flows that are solely payments of principal and interest on the principal amount outstanding.

In this case, 7% Teleco bonds are financial assets for Safe Boards Ltd. and meets the above criteria. Hence, can be measured at amortised cost.

The bonds should be reported in the statement of financial position at 30 June 2023 at 500 * 100 = $50,000.

Answer to Part b :

If the bonds were acquired for speculative purposes, then it should be reported in the statement of financial position at 30 June 2023 at 500 * 105 = $ 52,500. Thus, a change in fair value (market value) is recognised because the bond is held for trading/ speculative purpose. This fair value change of 500 * (105 - 100) = $ 2,500 should be recognised in Statement of Profit and Loss (FVTPL). This is because as per IFRS 9, a financial asset is classified through PL (FVTPL) if they are not held in one of the two business models mentioned above. (i.e not meant for holding and not meant for periodic interest and principal repayment - Which means held for speculation)